OpenStudy (nudeoua):

Solve 2a − 5d = 30 for d. Show your work How do i do it? step by step ?

OpenStudy (princeharryyy):

2a-5d+5d=30+5d 2a=30+5d 2a-30=30+5d+30 2a-30=5d d=(2a-30)/5

OpenStudy (wolf1728):

2a − 5d = 30 Subtracting 2a from both sides: -5d = 30 -2a Dividing both sides by -5: d = (30 -2a) / -5

imqwerty (imqwerty):

(:

OpenStudy (nudeoua):

Thanks!

OpenStudy (wolf1728):

u r welcome

imqwerty (imqwerty):

np B)

OpenStudy (wolf1728):

By the way both answers are correct. Let's suppose a = 25 and d=4 prince haryyy's answer d=(2a-30)/5 (2a-30)/5 = (50-30)/5 = 4 My answer d = (30 -2a) / -5 (30 -2a) / -5 = (30 -50) / -5 =4 So, both answers are the same.